OpenStudy (anonymous):

Which ratio is proportional to ? 8/18

OpenStudy (lgbasallote):

divide or multiply by a common factor

OpenStudy (anonymous):

so what would the answer be?

sam (.sam.):

\[\frac{8}{18}\div\frac{2}{2}=\frac{4}{9}\]

OpenStudy (lgbasallote):

a lot....4/9, 16/36, etc